# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ Appendix C Proofs in Section 2
For any $\bm{x},\bm{y}\in\mathbb{DH}^{n}$, we have
$\big\|\Pi(\bm{x}+\Pi(\bm{x})+\bm{y})-\Pi(\bm{x})\big\|_{2}\leq 2\|\bm{y}\|_{2}.$
@@
$\big\|\Pi(\bm{p}+\bm{q})-\bm{r}\big\|_{2}\leq 2\|\bm{q}-\bm{r}\|_{2}+3\big\|\Pi(\bm{p})-\bm{r}\big\|_{2}.$ (D.1)
By letting $p=cr$ for $c\in\mathbb{R}_{++}$ in (D.1), and forcing $c$ to zero, we can find that Lemma D.3 becomes
$\|\Pi(\bm{q})-\bm{r}\|_{2}\leq 2\|\bm{q}-\bm{r}\|_{2},$
which coincides with the result in Lemma 2.5. Lastly, to give a full picture of the proof for Theorem 3.2, we introduce the following lemma that demonstrates geometric properties of $\mathrm{UDQ}^{n}$.
@@
$z_{k}\in\operatorname{argmin}_{z\in\mathrm{UDQ}}\|\bm{x}^{k}-\hat{\bm{x}}z\|_{2},$ (D.2)
@@
The right hand side of (D.2) is the squared distance between $\bm{x}^{k}$ and $\hat{\bm{x}}$ after alignment, thus Lemma D.4 can be interpreted as an error bound condition. This lemma implies that sequence $\{\bm{x}^{(k)}\}_{k\geq 0}$ generated by GPM also satisfies an error bound condition, which plays an important role in proving Theorem 3.2 by bounding estimation error with distance terms. Additionally, as a degenerated case for SE(3), elements in $\mathrm{SO}(3)$ represented by unit dual quaternions with zero infinitesimal part should also satisfy (D.2). This extension on $\mathrm{SO}(3)$ meets the error bound attained for $\mathrm{SO}(d)$ in Liu et al. (2023), but (D.2) is tighter for $\mathrm{SO}(3)$ case since the equality always holds.
@@
$\operatorname{d}(\bm{x}^{k+1},\hat{\bm{x}})\leq\left(\frac{2}{\sqrt{n}}\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})+\frac{4}{n}\|\bm{\Delta}\|_{\mathrm{op}}\right)\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2}.$ (D.3)
Next, we prove by induction that, for all $k=0,1,2,\dots$,
$\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}\quad\text{and}\quad\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{18}.$ (D.4)
For the base case $k=0$, the claim follows immediately from Proposition 2.4 with the assumption $\|\bm{\Delta}\|_{\mathrm{op},st}\leq\frac{n}{350}$, $\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}\leq\frac{n}{300}$. Assume next that (D.4) holds for some $k\geq 0$. It remains to show that $\operatorname{d}_{st}(\bm{x}^{k+1},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}$ and $\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{18}$ . Combining the inductive hypothesis with again the noise assumptions, the standard part for right-hand side of inequality (D.3) and $\|\bm{\Delta}\hat{\bm{x}}\|_{2}\leq\sqrt{n}\|\bm{\Delta}\|_{\mathrm{op}}$ yield
$\left(\frac{2}{\sqrt{n}}\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})+\frac{4}{n}\|\bm{\Delta}\|_{\mathrm{op},st}\right)\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{\sqrt{n}}\|\bm{\Delta}\|_{\mathrm{op},st}\leq\frac{16}{175}\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})+\frac{6}{175}\sqrt{n}<\frac{\sqrt{n}}{25},$
and the dual part for right-hand side of (D.3) is
$\frac{4}{\sqrt{n}}\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{\sqrt{n}}\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}+\frac{4}{n}\big(\|\bm{\Delta}\|_{\mathrm{op},st}\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})+\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}\big)$
$\leq$ $\frac{4}{\sqrt{n}}\cdot\frac{n}{450}+\frac{12}{\sqrt{n}}\cdot\frac{n}{300}+\frac{4}{n}\left(\frac{n\sqrt{n}}{6300}+\frac{n\sqrt{n}}{7500}\right)<\frac{\sqrt{n}}{18},$
both of which close the induction and proves (D.4) for all $k\geq 0$.
Finally, using (D.3) again together with the fact that $\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}$ and $\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{18}$, we obtain the linear recursion for standard part:
$\operatorname{d}_{st}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{16}{175}\operatorname{d}_{st}(\bm{x}^{k-1},\hat{\bm{x}})+\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}.$
@@
$\leq\left(\frac{16}{175}\right)^{k}\operatorname{d}_{st}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}$
$<\left(\frac{1}{10}\right)^{k}\operatorname{d}_{st}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}.$
For the dual part, the recursion relation is
$\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{28}{225}\operatorname{d}_{st}(\bm{x}^{k-1},\hat{\bm{x}})+\frac{16}{175}\operatorname{d}_{\mathcal{I}}(\bm{x}^{k-1},\hat{\bm{x}})+\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}.$
Unrolling it gives
@@
$\leq\left(\frac{16}{175}\right)^{k}\operatorname{d}_{\mathcal{I}}(\bm{x}^{0},\hat{\bm{x}})+\left(1+\frac{16}{175}+\left(\frac{16}{175}\right)^{2}+\cdots\right)\left(\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}+\frac{5}{3n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}\right)$
@@
$<\left(\frac{1}{10}\right)^{k}\operatorname{d}_{\mathcal{I}}(\bm{x}^{0},\hat{\bm{x}})+k\left(\frac{1}{10}\right)^{k-1}\frac{1}{8}\operatorname{d}_{st}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}+\frac{875}{477n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}},$
which completes the proof. ∎
@@
$\mathcal{N}(ax)=\mathcal{N}(x),\quad\forall\,x\in\mathbb{DH}\setminus\{0\},\ a\in\mathbb{R}_{++}.$ (D.5)
If $x_{\mathrm{st}}\neq 0$, then by the definition of $\mathcal{N}(\cdot)$ and $|ax_{st}|=a|x_{st}|$, the standard part satisfies
$\frac{(ax)_{st}}{|(ax)_{\mathrm{st}}|}=\frac{ax_{\mathrm{st}}}{a|x_{\mathrm{st}}|}=\frac{x_{\mathrm{st}}}{|x_{\mathrm{st}}|},$
@@
$\frac{ax_{\mathcal{I}}}{|ax_{st}|}-\frac{ax_{st}}{|ax_{st}|}\mathrm{sc}\!\left(\frac{(ax_{\mathrm{st}})^{*}(ax_{\mathcal{I}})}{|ax_{\mathrm{st}}|\,|ax_{\mathrm{st}}|}\right)=\frac{x_{\mathcal{I}}}{|x_{st}|}-\frac{x_{st}}{|x_{st}|}\mathrm{sc}\!\left(\frac{x_{\mathrm{st}}^{*}x_{\mathcal{I}}}{|x_{\mathrm{st}}|\,|x_{\mathrm{st}}|}\right),$
which establishes (D.5) in this case.
@@
and the feasibility condition for $u_{\mathcal{I}}$ is invariant under positive scaling. Hence, (D.5) holds in this case as well.
Next, we prove that the projection $\Pi(\cdot)$ is scale-invariant under positive real scaling factor for any $\bm{y}\in\mathbb{DH}^{n}$. If $\bm{y}=\bm{0}$, then $a\bm{y}=\bm{0}$, and by definition, $\Pi(a\bm{y})=\bm{1}^{n}=\Pi(\bm{y})$. Otherwise, let $\tilde{\bm{y}}:=\Pi(\bm{y})$ and $\widetilde{\bm{y}}_{a}:=\Pi(a\bm{y})$. For each index $i$, if $y_{i}\neq 0$, then $(a\bm{y})_{i}\neq 0$, and by (D.5),
@@
If $y_{i}=0$, the definition uses $\mathcal{N}(\bm{e}^{*}\bm{y})$ for any $\bm{e}$ with $\bm{e}^{*}\bm{y}\neq 0$. Since $a>0$, the same $\bm{e}$ satisfies $\bm{e}^{*}(a\bm{y})=a(\bm{e}^{*}\bm{y})\neq 0$, and again by (D.5),
@@
Notation used in this proof: For given $\bm{u}=(u_{1},u_{2},\ldots,u_{n}),\bm{v}=(v_{1},v_{2},\ldots,v_{n})\in\mathbb{DH}^{n}$, with $u_{i}=u_{i,st}+u_{i,\mathcal{I}}\epsilon,v_{i}=v_{i,st}+v_{i,\mathcal{I}}\epsilon$. Denote $\langle\bm{u},\bm{v}\rangle$ as $(\bm{u}^{*}\bm{v}+\bm{v}^{*}\bm{u})/2$. Thus we have
$\langle\bm{u},\bm{v}\rangle=\sum_{i=1}^{n}\mathrm{sc}(u_{i,st}v_{i,st}^{*})+\sum_{i=1}^{n}\bigl(\mathrm{sc}(u_{i,st}v_{i,\mathcal{I}}^{*})+\mathrm{sc}(v_{i,st}u_{i,\mathcal{I}}^{*})\bigr)\epsilon.$
@@
$\|\bm{x}+\Pi(\bm{x})+\bm{y}-\bm{z}\|_{2}^{2}\leq\|\bm{x}+\Pi(\bm{x})+\bm{y}-\Pi(\bm{x})\|_{2}^{2}.$ (D.6)
Using $\|\bm{u}-\bm{v}\|_{2}^{2}=\|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}-2\langle\bm{u},\bm{v}\rangle$ for any $\bm{u},\bm{v}$ and $\|\bm{z}\|_{2}^{2}=\|\Pi(\bm{x})\|_{2}^{2}=n$ (since $\bm{z},\Pi(\bm{x})\in\mathrm{UDQ}^{n}$), inequality (D.6) simplifies to
$\langle\bm{y},\bm{z}-\Pi(\bm{x})\rangle\geq\langle\bm{x}+\Pi(\bm{x}),\Pi(\bm{x})-\bm{z}\rangle.$ (D.7)
@@
$\langle\bm{x},\Pi(\bm{x})-\bm{z}\rangle\geq 0.$ (D.8)
Combining (D.7) and (D.8), we obtain
@@
$\langle\bm{y},\bm{z}-\Pi(\bm{x})\rangle\geq\frac{1}{2}\|\Pi(\bm{x})-\bm{z}\|_{2}^{2}\geq 0.$ (D.9)
@@
Together with (D.9), this yields
@@
which proves the lemma.
@@
Combining the above three displays gives (D.1).

### v2

@@ Appendix C Proofs in Section 2
For any $x,y\in{\mathbb{DH}^{n}}$, we have
$\left\|\Pi(\bm{x}+\Pi(\bm{x})+\bm{y})-\Pi(\bm{x})\right\|_{2}\leq 2\|\bm{y}\|_{2}.$
@@
$\big\|\Pi(\bm{p}+\bm{q})-\bm{r}\big\|_{2}\leq 2\|\bm{q}-\bm{r}\|_{2}+3\big\|\Pi(\bm{p})-\bm{r}\big\|_{2}.$ (20)
By letting $\bm{p}=c\bm{r}$ for $c\in\mathbb{R}_{++}$ in (20), and forcing $c$ to zero, we can find that Lemma D.3 becomes
$\|\Pi(\bm{q})-\bm{r}\|_{2}\leq 2\|\bm{q}-\bm{r}\|_{2},$ (21)
which coincides with the result in Lemma 2.5.
Lemma D.4.
Let $r\in{\rm UDQ}^{n}$ and $h\in\mathbb{DH}^{n}$. Assume that
$\min_{1\leq i\leq n}|(r_{i}+h_{i})_{\rm st}|\geq\frac{1}{2},$ (22)
then
$\left\|\bigl(\Pi(\bm{r}+\bm{h})-\bm{r}\bigr)_{\mathcal{I}}\right\|_{2}\leq 2\|\bm{h}_{\mathcal{I}}\|_{2}+6\max_{1\leq i\leq n}|(r_{i})_{\mathcal{I}}|\,\|\bm{h}_{\rm st}\|_{2}.$ (23)
To give a full picture of the proof for Theorem 3.2, we introduce the following lemma that demonstrates geometric properties of $\mathrm{UDQ}^{n}$.
@@
$z_{k}\in\operatorname{argmin}_{z\in\mathrm{UDQ}}\|\bm{x}^{k}-\hat{\bm{x}}z\|_{2},$ (24)
@@
The right hand side of (24) is the squared distance between $\bm{x}^{k}$ and $\hat{\bm{x}}$ after alignment, thus Lemma D.5 can be interpreted as an error bound condition. This lemma implies that sequence $\{\bm{x}^{(k)}\}_{k\geq 0}$ generated by DQGPM also satisfies an error bound condition, which plays an important role in proving Theorem 3.2 by bounding estimation error with distance terms. Additionally, as a degenerated case for SE(3), elements in $\mathrm{SO}(3)$ represented by unit dual quaternions with zero infinitesimal part should also satisfy (24). This extension on $\mathrm{SO}(3)$ meets the error bound attained for $\mathrm{SO}(d)$ in Liu et al. (2023), but (24) is tighter for $\mathrm{SO}(3)$ case since the equality always holds.
Lastly, the below theorem gives a recursive relation for the dual part of the sequence $\{\bm{x}^{(k)}\}_{k\geq 0}$ generated by DQGPM.
Lemma D.6.
Under Assumption of Theorem 3.2, the aligned dual residual
$\delta_{\mathcal{I}}^{k}=\left\|\bigl(\bm{x}^{k}-\bm{r}^{k}\bigr)_{\mathcal{I}}\right\|_{2},\quad\text{where}\ \ \bm{r}^{k}=\hat{\bm{x}}z_{k},$
satisfies
$\delta_{\mathcal{I}}^{k+1}$ $\leq\left(\frac{2}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{2}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)\delta_{\mathcal{I}}^{k}$ (25)
[…]
$\quad+\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
@@
$\operatorname{d}(\bm{x}^{k+1},\hat{\bm{x}})\leq\left(\frac{2}{\sqrt{n}}\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})+\frac{4}{n}\|\bm{\Delta}\|_{\mathrm{op}}\right)\operatorname{d}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2}.$ (26)
First, we analyse the standard part. Recall that inequality (26) is in the lexicographic order of dual numbers. Hence it directly implies the following recursive relation for the standard component:
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{k+1},\hat{\bm{x}})\leq\left(\frac{2}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{4}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
At this stage, we only use the standard component of inequality (26); the infinitesimal component will be controlled separately. We prove by induction that, for all $k=0,1,2,\dots$,
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}.$ (27)
For the base case $k=0$, the claim follows immediately from Theorem 2.8 with the assumption $\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\leq\frac{n}{350}$. Assume next that (27) holds for some $k\geq 0$. It remains to show that $\operatorname{d}_{\mathrm{st}}(\bm{x}^{k+1},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}$. Combining the inductive hypothesis with again the noise assumptions, the standard part for right-hand side of inequality (26) and $\|\bm{\Delta}\hat{\bm{x}}\|_{2}\leq\sqrt{n}\|\bm{\Delta}\|_{\mathrm{op}}$ yield
$\left(\frac{2}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{4}{n}\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\right)\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{12}{\sqrt{n}}\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\leq\frac{16}{175}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{6}{175}\sqrt{n}<\frac{\sqrt{n}}{25},$
which closes the induction and proves (27) for all $k\geq 0$.
Using (26) again together with the fact that $\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{\sqrt{n}}{25}$, we obtain the linear recursion for standard part:
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\leq\frac{16}{175}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k-1},\hat{\bm{x}})+\frac{12}{n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}.$
@@
$\leq\left(\frac{16}{175}\right)^{k}\operatorname{d}_{\mathrm{st}}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}$
$<\left(\frac{1}{10}\right)^{k}\operatorname{d}_{\mathrm{st}}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,{st}}.$
Second, we control the dual coefficient. Unlike the standard component, this part is not obtained by taking the infinitesimal component of inequality (26). Instead, we use the aligned dual residual. Let
$\bm{u}=\bm{x}^{k}-\hat{\bm{x}}z_{k}=\bm{u}_{\rm st}+\epsilon\bm{u}_{\mathcal{I}}.$
If $\bm{u}_{\rm st}\neq 0$, then by the definition of the dual-valued norm, we know that
$(\|\bm{u}\|_{2})_{\mathcal{I}}=\frac{\sum_{i}{\rm sc}\bigl((\bm{u}_{i})_{\rm st}^{*}(\bm{u}_{i})_{\mathcal{I}}\bigr)}{\|\bm{u}_{\rm st}\|_{2}}.$
By the ordinary Cauchy–Schwarz inequality on $\mathbb{H}^{n}\simeq\mathbb{R}^{4n}$, we have
$\left|(\|\bm{u}\|_{2})_{\mathcal{I}}\right|\leq\|\bm{u}_{\mathcal{I}}\|_{2}.$
If $\bm{u}_{\rm st}=0$, then
$\|\bm{u}\|_{2}=\epsilon\|\bm{u}_{\mathcal{I}}\|_{2},$
@@
$\left|\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\right|\leq\left\|\bigl(\bm{x}^{k}-\hat{\bm{x}}z_{k}\bigr)_{\mathcal{I}}\right\|_{2}=\delta_{\mathcal{I}}^{k}.$
It remains to bound $\delta_{\mathcal{I}}^{k}$. By Lemma D.6, we have
@@
$\quad+\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
The standard-part basin bound and the noise assumption imply that the coefficient of $\delta_{\mathcal{I}}^{k}$ is strictly contractive:
$\frac{2}{\sqrt{n}}\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})+\frac{2}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}\leq\frac{2}{25}+\frac{2}{350}=\frac{3}{35}<\frac{1}{10}.$
Moreover, by (27),
[…]
Thus the one-step bound simplifies to
[…]
$\quad+\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$
Unrolling this recursion gives, for every $k\geq 1$,
[…]
$\quad+\left[\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}\right]\sum_{\ell=0}^{k-1}\left(\frac{1}{10}\right)^{k-1-\ell}.$
We now substitute the standard-part estimate. From (6), we obtain
$\delta_{\mathcal{I}}^{k}$ $\leq\left(\frac{1}{10}\right)^{k}\delta^{0}_{\mathcal{I}}+\left(\frac{2}{n}\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+\frac{6B+2\gamma}{n}\|\bm{\Delta}\|_{{\rm op},{\rm st}}+\frac{4B+\gamma}{25}\right)\left[\frac{\sqrt{n}}{25}k\left(\frac{1}{10}\right)^{k-1}+\frac{7000}{477n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}\right]$
$\quad+\frac{10}{9}\left[\frac{2}{\sqrt{n}}\left(\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}+B\|\bm{\Delta}\|_{{\rm op},{\rm st}}\right)+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}\right].$
This proves (10) combining the assumptions on noise $\|\bm{\Delta}\|_{{\rm op},{\rm st}}$ and $\|\bm{\Delta}\|_{{\rm op},\mathcal{I}}$. Together with $|\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})|\leq\delta_{\mathcal{I}}^{k}$, it also gives the desired bound on the infinitesimal coefficient of the dual-valued distance. ∎
@@
$\mathcal{N}(ax)=\mathcal{N}(x),\quad\forall\,x\in\mathbb{DH}\setminus\{0\},\ a\in\mathbb{R}_{++}.$ (28)
If $x_{\mathrm{st}}\neq 0$, then by the definition of $\mathcal{N}(\cdot)$ and $|ax_{\mathrm{st}}|=a|x_{\mathrm{st}}|$, the standard part satisfies
$\frac{(ax)_{\mathrm{st}}}{|(ax)_{\mathrm{st}}|}=\frac{ax_{\mathrm{st}}}{a|x_{\mathrm{st}}|}=\frac{x_{\mathrm{st}}}{|x_{\mathrm{st}}|},$
@@
$\frac{ax_{\mathcal{I}}}{|ax_{\mathrm{st}}|}-\frac{ax_{\mathrm{st}}}{|ax_{\mathrm{st}}|}\mathrm{sc}\!\left(\frac{(ax_{\mathrm{st}})^{*}(ax_{\mathcal{I}})}{|ax_{\mathrm{st}}|\,|ax_{\mathrm{st}}|}\right)=\frac{x_{\mathcal{I}}}{|x_{\mathrm{st}}|}-\frac{x_{\mathrm{st}}}{|x_{\mathrm{st}}|}\mathrm{sc}\!\left(\frac{x_{\mathrm{st}}^{*}x_{\mathcal{I}}}{|x_{\mathrm{st}}|\,|x_{\mathrm{st}}|}\right),$
which establishes (28) in this case.
@@
and the feasibility condition for $u_{\mathcal{I}}$ is invariant under positive scaling. Hence, (28) holds in this case as well.
Next, we prove that the projection $\Pi(\cdot)$ is scale-invariant under positive real scaling factor for any $\bm{y}\in\mathbb{DH}^{n}$. If $\bm{y}=\bm{0}$, then $a\bm{y}=\bm{0}$, and by definition, $\Pi(a\bm{y})=\bm{1}=\Pi(\bm{y})$. Otherwise, let $\tilde{\bm{y}}:=\Pi(\bm{y})$ and $\widetilde{\bm{y}}_{a}:=\Pi(a\bm{y})$. For each index $i$, if $y_{i}\neq 0$, then $(a\bm{y})_{i}\neq 0$, and by (28),
@@
If $y_{i}=0$, the definition uses $\mathcal{N}(\bm{e}^{*}\bm{y})$ for any $\bm{e}$ with $\bm{e}^{*}\bm{y}\neq 0$. Since $a>0$, the same $\bm{e}$ satisfies $\bm{e}^{*}(a\bm{y})=a(\bm{e}^{*}\bm{y})\neq 0$, and again by (28),
@@
Proof.
Notation used in this proof: For given $\bm{u}=(u_{1},u_{2},\ldots,u_{n}),\bm{v}=(v_{1},v_{2},\ldots,v_{n})\in\mathbb{DH}^{n}$, with $u_{i}=u_{i,\mathrm{st}}+u_{i,\mathcal{I}}\epsilon,v_{i}=v_{i,\mathrm{st}}+v_{i,\mathcal{I}}\epsilon$. Denote $\langle\bm{u},\bm{v}\rangle$ as $(\bm{u}^{*}\bm{v}+\bm{v}^{*}\bm{u})/2$. Thus we have
$\langle\bm{u},\bm{v}\rangle=\sum_{i=1}^{n}\mathrm{sc}(u_{i,\mathrm{st}}v_{i,\mathrm{st}}^{*})+\sum_{i=1}^{n}\bigl(\mathrm{sc}(u_{i,\mathrm{st}}v_{i,\mathcal{I}}^{*})+\mathrm{sc}(v_{i,\mathrm{st}}u_{i,\mathcal{I}}^{*})\bigr)\epsilon.$
@@
$\|\bm{x}+\Pi(\bm{x})+\bm{y}-\bm{z}\|_{2}^{2}\leq\|\bm{x}+\Pi(\bm{x})+\bm{y}-\Pi(\bm{x})\|_{2}^{2}.$ (29)
Using $\|\bm{u}-\bm{v}\|_{2}^{2}=\|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}-2\langle\bm{u},\bm{v}\rangle$ for any $\bm{u},\bm{v}$ and $\|\bm{z}\|_{2}^{2}=\|\Pi(\bm{x})\|_{2}^{2}=n$ (since $\bm{z},\Pi(\bm{x})\in\mathrm{UDQ}^{n}$), inequality (29) simplifies to
$\langle\bm{y},\bm{z}-\Pi(\bm{x})\rangle\geq\langle\bm{x}+\Pi(\bm{x}),\Pi(\bm{x})-\bm{z}\rangle.$ (30)
@@
$\langle\bm{x},\Pi(\bm{x})-\bm{z}\rangle\geq 0.$ (31)
Combining (30) and (31), we obtain
@@
$\langle\bm{y},\bm{z}-\Pi(\bm{x})\rangle\geq\frac{1}{2}\|\Pi(\bm{x})-\bm{z}\|_{2}^{2}\geq 0.$ (32)
@@
Together with (32), this yields
@@
which proves the lemma. ∎
@@
Combining the above three displays gives (20).